Apex Surgery Center performs knee and joint surgeries Monday through Saturday throughout the year. Apex Surgery Center has the following material costs for two items used in its operating room. Standard costs for one surgery include Item A: 5 pieces at $50 each Item B: 10 pieces at $75 each During September the following data apply to the Center Surgeries completed = 1,500 Item A purchased and used: 8,000 pieces at $40 per piece Item B purchased and used: 14,000 pieces at $76 per piece Compute Materials price and efficiency variances for Item A. Compute Materials price and efficiency variances for Item B.
Answer #1

1) Material price variance for Product A: (Actual Quantity Purchased * Actual Unit Cost) - (Actual Quantity Purchased * Standard Unit Cost)

= 320,000 - 400,000 = 80,000 Favorable

 

Efficiency variance for Product A: (Actual unit usage - Standard unit usage) * Standard cost per unit

= (8000-7500) * 50 = 25,000 Unfavorable

 

Material price variance for Product B: (Actual Quantity Purchased * Actual Unit Cost) - (Actual Quantity Purchased * Standard Unit Cost)

= 1,064,000 - 1,050,000 = 14,000 Unfavorable

 

Efficiency variance for Product B: (Actual unit usage - Standard unit usage) * Standard cost per unit

= (14,000 - 15,000) * 75 = 75,000 Favorable

 

Working:

Units

Actual cost Per unit

Total cost

A

8000

40

320,000

B

14000

76

1,064,000

1,384,000

Units

Standard cost Per unit

Total cost

A

8000

50

400,000

B

14000

75

1,050,000

1,450,000

Actual units

Standard units

Standard cost Per unit

A

8000

7500

50

B

14000

15000

75